# Mon CHÉRI: Mitigating Uninitialized Memory Access with Conditional Capabilities

### v2

@@ 1 Introduction
[…]
@@
In this paper, we extend the CHERI capability model to express memory-access policies that eliminate undefined behavior associated with uninitialized memory. We introduce conditional permissions (CPs) to capability-based addressing. CPs can express memory-access policies that take previous operations on memory into account. This enables conditional capabilities with fine-grained policies that satisfy different instances of memory-safety objectives at different granularities, such as “no reads of memory which has not been the subject of at least one write” (Write-before-Read) or “this memory can be written to only once” (Write-Once). We describe these CPs in § 3. CPs are enforced by introducing the notion of operation-specific bounds to capability-based addressing:
  • We introduce conditional permissions and conditional capabilities for capability-based addressing (§ 3).
  • We integrate conditional capabilities to CHERI-RISC-V in prototypes based on the QEMU full-system emulator and an FPGA softcore based on Flute64Cheri IP (§ 4).
  • We add support for Write-before-Read conditional permissions to the CHERI-enabled Clang/LLVM compiler (§ 4.3) and embedded memory allocators (§ 4.4).
  • We evaluate CPs using $>1000$ NIST Juliet test suite test cases for uninitialized variables and using EEMBC CoreMark performance benchmarks (§ 5).
@@ 2.2 The CHERI Capability Architecture
Figure 1 shows the in-memory representation of a CHERI capability. Each capability is double the width of the native integer pointer type of the baseline architecture: 128 bits on 64-bit platforms and 64 bits on 32-bit platforms. One additional bit, the validity tag ➀, is stored separate from the capability and is protects its integrity: any manipulation of the capability in-memory by non-capability-aware instructions invalidates its tag. Capability-aware instructions maintain the tag as long as certain architectural invariants are met. This prevents direct in-memory manipulations and injection of arbitrary data as capabilities. The permissions ➁ control how the capability can be used and consists of the permissions described in § 2.1. The object type ➂ allows capabilities to be temporarily “sealed”, which renders the capability unusable until it is “unsealed” by a special instruction. Sealing is used by CHERI to implement opaque pointer types and fine-grained in-process isolation. The bounds ➃ describe a lower and upper bound relative to the baseline architecture address ➄, which limits the portion of address space the capability is able to access. To reduce the in-memory footprint of capabilities, the bounds are stored in a compressed format [19] with both bounds in 28 bits (for a 64-bit address), loosing precision as the object size increases.
@@ 3 Conditional Capability Design
TABLE I: Conditional permissions types for conditional capabilities
Conditional permissions Description
@@ 3.1 Challenges
Integrating CPs into capability-based addressing presents several challenges. A key issue is managing the tracking of uninitialized data. Previous approaches [25, 26, 27] assume that data between the lower bound and address pointed to by the capability is initialized, while uninitialized data falls between the address and the upper bound. However, this approach has limitations, as discussed in § 6. To address this, we propose operation-specific bounds (OBs), which precisely track writes. OBs must be updated when a program writes to a region of memory to reflect the new bounds. Otherwise, stale or misaligned bounds can lead to false positives.
@@
To store and update OBs efficiently, we propose leveraging unused bits in the baseline architecture address (Figure 1). This minimizes changes to the underlying hardware but makes significant compression of the OB necessary. Further details on the hardware modifications are provided in § 3.3 and § 4.
@@
To be practical, CPs must integrate seamlessly with existing capability architectures, such as CHERI, that are gaining industry adoption. As a case study, we integrate CPs to CHERI while ensuring full compatibility with the underlying RISC-V ISA. In § 4, we describe the integration process and our hardware extension, “Mon CHÉRI”, which introduces only minor modifications to the CHERI capability representation (§ 4.2).
@@
CPs add, similar to capability sealing (§ 2.2), controlled non-monotonicity to the CHERI design. They allow temporary suspension of permissions, which are regained once the associated condition is met. Our design (§ 3.2) guarantees that CPs do not to elevate privileges beyond the original capabilities.
@@
Capability linearity refers to the consistency and integrity of OBs across a program’s execution. To maintain this, we propose a new compiler pass to ensure that capabilities are properly updated as the program executes. This pass is implemented in the LLVM-based Mon CHÉRI toolchain, as detailed in § 3.3 and § 4.3.
@@
A concern with any memory access control mechanism is the potential for performance degradation. For Mon CHÉRI, we ensure that checks involved in monitoring memory accesses do not add significant latency or computational overhead during normal program execution. In § 4.1, we explain how the OB checks are carefully designed to minimize additional cycles per instruction.
@@ 3.2 Conditional Capabilities
$\mathcal{P}=\begin{cases}\mathcal{P^{U}}&if\ \mathcal{P^{C}}=\emptyset\\ \mathcal{P^{U}}\land\left(\mathcal{X}\implies\mathcal{P^{C}}\right)&if\ \mathcal{P^{C}}\neq\emptyset\\ \end{cases}$
@@
Figure 2 illustrates an example use case for our conditional capabilities: Write-before-Read. In its initial state (Figure 2.a), the capability refers to a writable memory area with upper and lower bounds. If a store occurs in the memory area (Figure 2.b), the OB increases by the size of the store operand on the hardware level.
An advantage of our design is that it allows capability hardware to enforce a variety of novel access-control permissions beyond conventional RWX. We identify the CPs and corresponding use cases in Table I. This set of CPs allows the definition of “single-use” code and data that enable policies that can be used to harden software against run-time attacks that aim to reuse existing program resource. Such attacks, e.g., code-reuse attacks such as return-oriented-programming [28], can use pre-existing sequences of instructions to craft malicious control flows that can be used to compromise program behavior. Code that a process executes only once, e.g., a fixed sequence of initialization instructions, could, using the Execute-Once CP be invoked once during program initialization and is subsequently rendered useless for code-reuse attacks later in the process’ lifetime.
@@ 3.3 High-Level System Architecture
Figure 3 depicts the high-level system architecture for our prototype conditional capability-enhanced LLVM compiler and CHERI processor. Inputs and process blocks in a darker color indicate the changes to a conventional CHERI-enabled compiler and processor architecture. For conciseness, in this description, we focus on Write-before-Read CPs.
In our prototype, conditional capabilities are exposed to developers through the three alternative usage models: A a Write-before-Read compiler option, B function-level annotations, and C variable-level annotations. Our current design of the compiler option A applies Write-before-Read to all stack variables, but we discuss possibilities for more intelligent heuristics in § 7. Function- and variable-level annotations are implemented as Clang function and variable attributes, respectively (shown in LABEL:lst:attributes), which gives developers control over which variables Write-before-Read is applied to.
The conditional capability-enhanced LLVM compiler (Figure 3(a)) can operate either on original source code or source code annotated by the developer using the aforementioned attributes. This differs from the conventional CHERI-enabled LLVM in three ways: ➀ the CP instrumentation transform pass that marks variable for instrumentation based on the compiler option or developer annotations, and emits intrinsics that interface with the conditional capability hardware to initialize operation-specific bounds for newly created capabilities, ➁ a store linearization transform pass that ensures that accesses to uninitialized objects in memory are performed consistently through a capability that tracks the operation bound, and ➂ support for the new conditional capability instructions in the CHERI RISC-V back end, allowing the compiler to interface with the conditional capability-enhanced CHERI processor.
The conditional capability-enhanced CHERI processor (Figure 3(a)) is modified to support the conditional capability instructions (➃, § 4.1), and its instruction pipeline is augmented with logic to encode and decode capabilities containing an additional operation bound (➄, § 3.2). Modified ➅ store and ➆ load logic updates and checks the operation bounds when executing store and load instructions, respectively. Finally, ➇ the pipeline performs writebacks of updated capabilities in register operands after the operations that update the operation bounds. Code capability checks for the Program Counter Capability (PCC) are done in a dedicated bounds check block ➈ (needed for Write-before-Execute, Write-before-Execute-Only , and Execute-Once), while all data CPs use a general-purpose bounds check block ➉.
@@ 4 Mon CHÉRI Implementation
In this section, we present Mon CHÉRI, our implementation of CPs for the CHERI-RISC-V architecture. We implemented two versions of Mon CHÉRI: 1. a Mon CHÉRI software model for the QEMU-system-CHERI128 full-system emulator, and 2. a Mon CHÉRI softcore based on the CHERI-Flute64 processor intellectual property (IP). The Mon CHÉRI ISA extension is described in § 4.1.
In addition, we extended the existing CHERI-LLVM toolchain [30] with support for the Mon CHÉRI ISA extension and Write-before-Read CP instrumentation for stack-allocated variables (§ 4.3). Finally, we added run-time Write-before-Read CP support for heap-allocated memory in the CheriFreeRTOS [31] and Two-Level Segregated Fit (TLSF) [32] memory allocators (§ 4.4).
@@ 4.1 Mon CHÉRI Extension for CHERI-RISC-V
The RISC-V ISA is extensible through portions of the instruction encoding space reserved for ISA extensions. We add a family of CSetOpBounds instructions to CHERI-RISC-V to the existing CHERIv9 [16] non-standard Xcheri extension in the custom-2/rv128 opcode space to allow setting an initial, or updating an already set, value for a capability’s operation bound. Our CSetOpBounds instructions require two operands: an existing capability and a length operand. When invoked, CSetOpBounds sets the operation bound to the range $\left[b,o\right]$ where $b$ is the “base” encoded as part of the conventional capability bound, and $o$ is the “operation top” encoded into the operation bound (see § 4.2). CSetOpBounds does not allow programs to increase the operation bounds, while they are allowed to invoke CSetOpBounds to decrease the operation bounds of a capability. The operational bounds set by CSetOpBounds are restricted to be within the original capability bounds.
@@
We assign new conditional permissions control bits from the CHERI capability representation ($p_{op}$ in Figure 4(b)). The CHERI-Flute64 IP reserves 12-bits for hardware-defined permissions, i.e., those authorizing load, store, etc., and 4-bits for software-defined permissions, which interpretation are left open by the CHERIv9 specification. For the purposes of prototyping CPs, we utilize the four available software-defined bits, treating them as a 4-bit integer that enumerates one of 16 possible, mutually exclusive permission states. Five of those states correspond to the CPs in table II, ten modes are left unused, and one mode, the zero value, corresponds to the default state in which CP enforcement is disabled. In this default state, protections such as Write-before-Read are inactive, allowing uninitialized memory access within the capability’s bounds. Invoking csetwbropbound on a capability enables Write-before-Read enforcement for that capability.
@@
[…]
@@
[…]
@@
To solve this, we implemented a bypass—a new data path within the pipeline—that forwards the updated OB from the memory access or writeback stage to the ALU stage. This bypass activates when the processor detects that the next instruction in the pipeline operates on the same conditional capability. This avoids the need to stall the processor or reorder instructions during compilation and is compatible with conventional RISC-V instruction ordering. Appendix § A.1 shows an example of a data hazard that is avoided by the bypass.
@@ 4.2 Adding Operation Bounds to Capabilities
Figure 4(a) shows the capability format for CHERI Concentrate. $B$ and $T$ encode the $b$ and $t$ bounds in one of two formats depending on the internal exponent ($I_{E}$) bit: if $I_{E}=1$ then an exponent ($E$) is stored in the lower three bits of $B$ and $T$ ($B_{E}$ and $T_{E}$) reducing their precision by three bits. $E$ determines the position at which $B$ and $T$ are inserted into $a$ to obtain $b$ and $t$. Otherwise ($I_{E}=0$, $E=0$) the full width of $b$ and $t$ are used. Their width is determined by an encoding parameter: mantissa width ($MW$) that determines the precision of the decoded bounds. The CHERI ISAv9 uses $MW=14$ for 128-bit capabilities. However, $t$ is further compressed by two bits as the top two bits of $t$ can be derived from the equation $T=B+L$ where the most significant bit of $L$ ($L_{msb}$) is known from the values of $I_{E}$ and $E$ and a carry bit is implied if $T\left[11:0\right]<B\left[11:0\right]$ since $t$ is known to be larger than $b$.
@@
Figure 4(b) shows the capability format for Mon CHÉRI. We encode the operation bound, operation top ($o$) in two fields in the most significant 16-bits of the cursor: $O\left[13:3\right]$ (11 bits) and $O_{E}$ (5 bits) which are freed by limiting $a_{top}=a\left[47:E+MW\right]$. When $I_{E}=0$, $o$ is stored identically to $b$ with its lowest three bits $O\left[2:0\right]$ derived from the most significant bits of $O_{E}$ ($O_{E}\left[4:2\right]$). When $I_{E}=1$, up to five bits from $O_{E}\left[E+2:0\right]$ are used to store the least significant bits of $o$. Our current implementation limits $O_{E}$ to five bits, limiting $o$ to an $I_{E}$ of at most $2$. We discuss methods to alleviate this limitation in § 7. Expanding $o$ to full 48 bits happens similarly to $b$ with its own correction $c_{o}$. Figure 6 in § A.2 illustrates the changes relative to CHERI Concentrate compression.

### v3

@@ 1 Introduction
➁Permissions
➂Object type
➃Bounds
➄Baseline architecture address
@@
In this paper, we extend the CHERI capability model to express memory-access policies that eliminate undefined behavior associated with uninitialized memory. We introduce conditional permissions (CPs) to capability-based addressing. CPs can express memory-access policies that take previous operations on memory into account. This enables conditional capabilities with fine-grained policies that satisfy different instances of memory-safety objectives at different granularities, such as “no reads of memory which has not been the subject of at least one write” (Write-before-Read) or “this memory can be written to only once” (Write-Once). We describe these CPs in §˜3. CPs are enforced by introducing the notion of operation-specific bounds to capability-based addressing:
  • We introduce conditional permissions and conditional capabilities for capability-based addressing (§˜3).
  • We integrate conditional capabilities to CHERI-RISC-V in prototypes based on the QEMU full-system emulator and an FPGA softcore based on Flute64Cheri IP (§˜4).
  • We add support for Write-before-Read conditional permissions to the CHERI-enabled Clang/LLVM compiler (§˜4.3) and embedded memory allocators (§˜4.4).
  • We evaluate CPs using >1000>1000 NIST Juliet test suite test cases for uninitialized variables and using EEMBC CoreMark performance benchmarks (§˜5).
@@ 2.2 The CHERI Capability Architecture
Figure˜1 shows the in-memory representation of a CHERI capability. Each capability is double the width of the native integer pointer type of the baseline architecture: 128 bits on 64-bit platforms and 64 bits on 32-bit platforms. One additional bit, the validity tag ➀, is stored separate from the capability and is protects its integrity: any manipulation of the capability in-memory by non-capability-aware instructions invalidates its tag. Capability-aware instructions maintain the tag as long as certain architectural invariants are met. This prevents direct in-memory manipulations and injection of arbitrary data as capabilities. The permissions ➁ control how the capability can be used and consists of the permissions described in §˜2.1. The object type ➂ allows capabilities to be temporarily “sealed”, which renders the capability unusable until it is “unsealed” by a special instruction. Sealing is used by CHERI to implement opaque pointer types and fine-grained in-process isolation. The bounds ➃ describe a lower and upper bound relative to the baseline architecture address ➄, which limits the portion of address space the capability is able to access. To reduce the in-memory footprint of capabilities, the bounds are stored in a compressed format [19] with both bounds in 28 bits (for a 64-bit address), loosing precision as the object size increases.
@@ 3 Conditional Capability Design
TABLE I: Conditional permission types for conditional capabilities
Conditional permission Description
@@ 3.1 Challenges
Integrating CPs into capability-based addressing presents several challenges. A key issue is managing the tracking of uninitialized data. Previous approaches [25, 26, 27] assume that data between the lower bound and address pointed to by the capability is initialized, while uninitialized data falls between the address and the upper bound. However, this approach has limitations, as discussed in §˜6. To address this, we propose operation-specific bounds (OBs), which precisely track writes. OBs must be updated when a program writes to a region of memory to reflect the new bounds. Otherwise, stale or misaligned bounds can lead to false positives.
@@
To store and update OBs efficiently, we propose leveraging unused bits in the baseline architecture address (Figure˜1). This minimizes changes to the underlying hardware but makes significant compression of the OB necessary. Further details on the hardware modifications are provided in §˜3.3 and §˜4.
@@
To be practical, CPs must integrate seamlessly with existing capability architectures, such as CHERI, that are gaining industry adoption. As a case study, we integrate CPs to CHERI while ensuring full compatibility with the underlying RISC-V ISA. In §˜4, we describe the integration process and our hardware extension, “Mon CHÉRI”, which introduces only minor modifications to the CHERI capability representation (§˜4.2).
@@
CPs add, similar to capability sealing (§˜2.2), controlled non-monotonicity to the CHERI design. They allow temporary suspension of permissions, which are regained once the associated condition is met. Our design (§˜3.2) guarantees that CPs do not to elevate privileges beyond the original capabilities.
@@
Capability linearity refers to the consistency and integrity of OBs across a program’s execution. To maintain this, we propose a new compiler pass to ensure that capabilities are properly updated as the program executes. This pass is implemented in the LLVM-based Mon CHÉRI toolchain, as detailed in §˜3.3 and §˜4.3.
@@
A concern with any memory access control mechanism is the potential for performance degradation. For Mon CHÉRI, we ensure that checks involved in monitoring memory accesses do not add significant latency or computational overhead during normal program execution. In §˜4.1, we explain how the OB checks are carefully designed to minimize additional cycles per instruction.
@@ 3.2 Conditional Capabilities
[…]
@@
Figure˜2 illustrates an example use case for our conditional capabilities: Write-before-Read. In its initial state (Figure˜2.a), the capability refers to a writable memory area with upper and lower bounds. If a store occurs in the memory area (Figure˜2.b), the OB increases by the size of the store operand on the hardware level.
An advantage of our design is that it allows capability hardware to enforce a variety of novel access-control permissions beyond conventional RWX. We identify the CPs and corresponding use cases in Table˜I. This set of CPs allows the definition of “single-use” code and data that enable policies that can be used to harden software against run-time attacks that aim to reuse existing program resource. Such attacks, e.g., code-reuse attacks such as return-oriented-programming [28], can use pre-existing sequences of instructions to craft malicious control flows that can be used to compromise program behavior. Code that a process executes only once, e.g., a fixed sequence of initialization instructions, could, using the Execute-Once CP be invoked once during program initialization and is subsequently rendered useless for code-reuse attacks later in the process’ lifetime.
@@ 3.3 High-Level System Architecture
Figure˜3 depicts the high-level system architecture for our prototype conditional capability-enhanced LLVM compiler and CHERI processor. Inputs and process blocks in a darker color indicate the changes to a conventional CHERI-enabled compiler and processor architecture. For conciseness, in this description, we focus on Write-before-Read CPs.
In our prototype, conditional capabilities are exposed to developers through the three alternative usage models: A a Write-before-Read compiler option, B function-level annotations, and C variable-level annotations. Our current design of the compiler option A applies Write-before-Read to all stack variables, but we discuss possibilities for more intelligent heuristics in §˜7. Function- and variable-level annotations are implemented as Clang function and variable attributes, respectively (shown in LABEL:lst:attributes), which gives developers control over which variables Write-before-Read is applied to.
The conditional capability-enhanced LLVM compiler (Figure˜3(a)) can operate either on original source code or source code annotated by the developer using the aforementioned attributes. This differs from the conventional CHERI-enabled LLVM in three ways: ➀ the CP instrumentation transform pass that marks variable for instrumentation based on the compiler option or developer annotations, and emits intrinsics that interface with the conditional capability hardware to initialize operation-specific bounds for newly created capabilities, ➁ a store linearization transform pass that ensures that accesses to uninitialized objects in memory are performed consistently through a capability that tracks the operation bound, and ➂ support for the new conditional capability instructions in the CHERI RISC-V back end, allowing the compiler to interface with the conditional capability-enhanced CHERI processor.
The conditional capability-enhanced CHERI processor (Figure˜3(a)) is modified to support the conditional capability instructions (➃, §˜4.1), and its instruction pipeline is augmented with logic to encode and decode capabilities containing an additional operation bound (➄, §˜3.2). Modified ➅ store and ➆ load logic updates and checks the operation bounds when executing store and load instructions, respectively. Finally, ➇ the pipeline performs writebacks of updated capabilities in register operands after the operations that update the operation bounds. Code capability checks for the Program Counter Capability (PCC) are done in a dedicated bounds check block ➈ (needed for Write-before-Execute, Write-before-Execute-Only , and Execute-Once), while all data CPs use a general-purpose bounds check block ➉.
@@ 4 Mon CHÉRI Implementation
In this section, we present Mon CHÉRI, our implementation of CPs for the CHERI-RISC-V architecture. We implemented two versions of Mon CHÉRI: 1. a Mon CHÉRI software model for the QEMU-system-CHERI128 full-system emulator, and 2. a Mon CHÉRI softcore based on the CHERI-Flute64 processor intellectual property (IP). The Mon CHÉRI ISA extension is described in §˜4.1.
In addition, we extended the existing CHERI-LLVM toolchain [30] with support for the Mon CHÉRI ISA extension and Write-before-Read CP instrumentation for stack-allocated variables (§˜4.3). Finally, we added run-time Write-before-Read CP support for heap-allocated memory in the CheriFreeRTOS [31] and Two-Level Segregated Fit (TLSF) [32] memory allocators (§˜4.4).
@@ 4.1 Mon CHÉRI Extension for CHERI-RISC-V
The RISC-V ISA is extensible through portions of the instruction encoding space reserved for ISA extensions. We add a family of CSetOpBounds instructions to CHERI-RISC-V to the existing CHERIv9 [16] non-standard Xcheri extension in the custom-2/rv128 opcode space to allow setting an initial, or updating an already set, value for a capability’s operation bound. Our CSetOpBounds instructions require two operands: an existing capability and a length operand. When invoked, CSetOpBounds sets the operation bound to the range $\left[b,o\right]$ where $b$ is the “base” encoded as part of the conventional capability bound, and $o$ is the “operation top” encoded into the operation bound (see §˜4.2). CSetOpBounds does not allow programs to increase the operation bounds, while they are allowed to invoke CSetOpBounds to decrease the operation bounds of a capability. The operational bounds set by CSetOpBounds are restricted to be within the original capability bounds.
@@
We assign new conditional permissions control bits from the CHERI capability representation ($p_{op}$ in Figure˜4(b)). The CHERI-Flute64 IP reserves 12-bits for hardware-defined permissions, i.e., those authorizing load, store, etc., and 4-bits for software-defined permissions, which interpretation are left open by the CHERIv9 specification. For the purposes of prototyping CPs, we utilize the four available software-defined bits, treating them as a 4-bit integer that enumerates one of 16 possible, mutually exclusive permission states. Five of those states correspond to the CPs in table˜II, ten modes are left unused, and one mode, the zero value, corresponds to the default state in which CP enforcement is disabled. In this default state, protections such as Write-before-Read are inactive, allowing uninitialized memory access within the capability’s bounds. Invoking csetwbropbound on a capability enables Write-before-Read enforcement for that capability.
@@
023131416172526274445464748596063
$p_{sw}$
$p_{hw}$
$f$
otype
$I_{E}$
$T\left[11:3\right]$
$T_{E}$
$B\left[13:3\right]$
$B_{E}$
$a$
Cursor
@@
023131416172526274445464748596063
$p_{op}$
$p_{hw}$
$f$
otype
$I_{E}$
$T\left[11:3\right]$
$T_{E}$
$B\left[13:3\right]$
$B_{E}$
$O\left[13:3\right]$
$O_{E}$
$a$
Cursor
@@
To solve this, we implemented a bypass—a new data path within the pipeline—that forwards the updated OB from the memory access or writeback stage to the ALU stage. This bypass activates when the processor detects that the next instruction in the pipeline operates on the same conditional capability. This avoids the need to stall the processor or reorder instructions during compilation and is compatible with conventional RISC-V instruction ordering. Appendix §˜A.1 shows an example of a data hazard that is avoided by the bypass.
@@ 4.2 Adding Operation Bounds to Capabilities
Figure˜4(a) shows the capability format for CHERI Concentrate. $B$ and $T$ encode the $b$ and $t$ bounds in one of two formats depending on the internal exponent ($I_{E}$) bit: if $I_{E}=1$ then an exponent ($E$) is stored in the lower three bits of $B$ and $T$ ($B_{E}$ and $T_{E}$) reducing their precision by three bits. $E$ determines the position at which $B$ and $T$ are inserted into $a$ to obtain $b$ and $t$. Otherwise ($I_{E}=0$, $E=0$) the full width of $b$ and $t$ are used. Their width is determined by an encoding parameter: mantissa width ($MW$) that determines the precision of the decoded bounds. The CHERI ISAv9 uses $MW=14$ for 128-bit capabilities. However, $t$ is further compressed by two bits as the top two bits of $t$ can be derived from the equation $T=B+L$ where the most significant bit of $L$ ($L_{msb}$) is known from the values of $I_{E}$ and $E$ and a carry bit is implied if $T\left[11:0\right]<B\left[11:0\right]$ since $t$ is known to be larger than $b$.
@@
Figure˜4(b) shows the capability format for Mon CHÉRI. We encode the operation bound, operation top ($o$) in two fields in the most significant 16-bits of the cursor: $O\left[13:3\right]$ (11 bits) and $O_{E}$ (5 bits) which are freed by limiting $a_{top}=a\left[47:E+MW\right]$. When $I_{E}=0$, $o$ is stored identically to $b$ with its lowest three bits $O\left[2:0\right]$ derived from the most significant bits of $O_{E}$ ($O_{E}\left[4:2\right]$). When $I_{E}=1$, up to five bits from $O_{E}\left[E+2:0\right]$ are used to store the least significant bits of $o$. Our current implementation limits $O_{E}$ to five bits, limiting $o$ to an $I_{E}$ of at most $2$. We discuss methods to alleviate this limitation in §˜7. Expanding $o$ to full 48 bits happens similarly to $b$ with its own correction $c_{o}$. Figure˜6 in §˜A.2 illustrates the changes relative to CHERI Concentrate compression.